Suppose NAD+ is unavailable because NADH cannot be oxidized due to a mutation in

the NADH dehydrogenase (Complex I). If FAD could substitute for all NAD+, how would

this affect the total ATP produced by the complete oxidation of glucose? How would it

affect the total ATP produced by the complete oxidation of a 20C fatty acid? (Give specific

numbers and show how you reached this number.) Based on what you know about redox

potentials, would this be possible?
Answer #1

ANSWER 1 :-

During Glycolysis process, 2 ATP molecules are directly produced along with two ATP being produced by means of Substrate level phosphorylation during the Kreb's cycle. So, direct production of ATP is 4.

There are total of 10 NADH2 molecules and 2 FADH2 molecules produced during glucose oxidation. If NADH2 is being substituted to FADH2 during ETS, there will be overall 12 molecules of NADH2 available.

Each NADH2 is able to produce 2 molecules of ATP and hence 12 molecules will be able to produce 24 molecules of ATP.

So, the net gain of ATP molecules will be ( 24 + 4 = 28 ) molecules of ATP which is reduced by 10 if NADH2 would have been utilised (38).

During Glycolysis, 2 molecules of NADH2 are produced. During the process where pyruvate is converted to Acetyl-CoA, 2 molecules of NADH2 are produced. Kreb's cycle for single molecule of Acetyl-CoA is responsible for producing 3 molecules of NADH2 and thus a total of 6 NADH2 for both the Acetyl-CoA entering the mitochondrial matrix.

ANSWER 2 :-

Lets consider the beta-oxidation of stearic acid. It is composed of 18 carbon atoms. Activation step will be associated with utilisation of 2 molecules of ATP. There will be 8 oxidation steps associated with the conversion to Acetyl CoA which will lead to production of 8 molecules of both NADH2 and FADH2.

Acetyl-CoA enters the Kreb's cycle where it will produce 1 ATP, 3 NADH2 and 1 FADH2 molecule. But, 9 Acetyl-CoA molecules will be produced. So, direct ATP production will be 9 ATP. 27 NADH2 and 9 FADH2 will be produced.

Each NADH2 produces 3 ATP and each FADH2 produced 2 ATP.

NADH2 are 35 which produces total of 105 ATP.

FADH2 are 17 which produces total of 34 ATP.

9 ATP are directly produced and 2 are used in activation step.

So, net gain of ATP is ( 105 + 34 + 9 - 2 = 146 ATP ).

Now, when NADH2 is substituted by FADH2, there will be total of 52 FADH2 which will be utilised which will produce total of 104 ATP molecules.

So, net gain of ATP will be ( 104 + 9 - 2 = 111 ATP ).

So, there is a decrease in ATP production.

Note :- Respected Sir, for any doubts please prefer communicating through comment section and please provide an upvote if the answer seems satisfactory.